# Continual Knowledge Adaptation for Reinforcement Learning

### v1

@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful paradigm in machine learning, enabling agents to learn optimal behaviors through interactions with dynamic environments kaelbling1996reinforcement ; dohare2024loss ; anonymous2025openworld . RL has achieved significant success in fields such as robotic control johannink2019residual ; meng2025preserving , embodied intelligence chen2022learning ; gupta2021embodied , and natural language processing hu2024dynamic ; guo2025deepseek ; hu2025dynamic . However, traditional RL usually assumes a static environment where tasks and data distributions remain fixed, intending to solve a single, well-defined problem malagonself . In contrast, real-world environments are typically non-stationary, requiring agents to continuously adapt to new tasks and evolving conditions. In light of this, Continual Reinforcement Learning (CRL) khetarpal2022towards has been introduced to enable agents to adapt and maintain performance across multiple tasks, facilitating more robust decision-making in dynamic environments gaya2023building ; abel2023definition ; wang2024comprehensive .
@@
Unfortunately, CRL still faces several challenges, which are as follows. 1) Cross-task conflict: In continual learning, tasks may share certain structures or knowledge while also having incompatible goals or constraints. This interplay complicates the direct reuse of previously learned information kumaran2016learning . 2) Catastrophic forgetting: When learning new tasks, the agent may overwrite or distort previously acquired knowledge, leading to a loss of performance on earlier tasks malagonself ; wang2024comprehensive .
@@
Recently, several attempts have been proposed to enable agents to continuously learn across multiple tasks while maintaining or enhancing their performance on previously acquired tasks zheng2025mastering ; hu2024continual . The existing methods can be broadly categorized into four main groups wang2024comprehensive . The regularization-based methods igltransient ; muppidi2024fast introduce regularization terms into the learning objective that penalize large updates to important parameters for previously learned tasks. The rehearsal-based methods daniels2022model ; zhang2023catastrophic ; zhang2023replayenhanced mitigate forgetting by storing previous experiences in memory and periodically replaying them during the learning of new tasks. By leveraging past experiences, these methods help reduce short-term biases and improve task performance across the sequence. The architecture-based methods ullah2021machine ; malagonself ; ben2022lifelong ; rusu2016progressive focus on learning a shared structure, such as modularity or composition, to facilitate continual learning. They reuse parts of previous solutions by forming abstract concepts or skills. The meta-learning based methods schmidhuber1998reinforcement ; finn2017model ; luo2022adapt ; clavera2018model improve the learning efficiency of agents by utilizing past successes and failures to refine their optimization processes. This creates an inductive bias that enhances sample efficiency and adaptability in acquiring new behaviors.
@@
Although there has been significant progress in existing methods, they still face several limitations as follows. Firstly, existing methods, such as PackNet mallya2018packnet , often fail to effectively address task dependencies and conflicts, which makes it challenging to transfer knowledge across tasks without interference or degradation in performance. Secondly, most methods still exhibit substantial performance degradation on previously learned tasks upon acquisition of new ones. This issue is particularly pronounced in methods that rely on regularization, as they may struggle to efficiently retain knowledge from earlier tasks as the task sequence lengthens. Lastly, many methods face scalability issues, especially as the number of tasks grows significantly, leading to increased memory and computational costs, such as CompoNet malagonself .
@@ 2 Related Work
Continual Reinforcement Learning (CRL) aims to enable agents to continuously learn and optimize strategies in dynamic environments, improving their ability to adapt to environmental changes and achieve goals more efficiently. The existing methods can be broadly categorized into four main types wang2024comprehensive : regularization-based methods, rehearsal-based methods, architecture-based methods, and meta-learning-based methods, each of which is detailed as:
@@
Regularization-based Methods. These methods igltransient ; muppidi2024fast add regularization terms to the training process to balance learning new and old tasks without storing models of old tasks. ITER igltransient periodically distills the current policy and value function into a newly initialized network during training and imitates the teacher network using a linear combination of loss terms to enhance model generalization. TRAC muppidi2024fast adaptively adjusts the regularization strength based on online convex optimization theory to prevent excessive weight drift, thereby reducing plasticity loss and enabling rapid adaptation to new distribution changes. Anand et al. anand2023prediction propose a method that decomposes the value function into permanent and transient components to address the stability-plasticity dilemma in continual reinforcement learning, enabling efficient adaptation and control in dynamic environments.
@@
Rehearsal-based Methods. These methods daniels2022model ; zhang2023catastrophic ; zhang2023replayenhanced ; fuknowledge utilize experience replay, generation replay, parameter replay, etc., to store past experiences and prevent catastrophic forgetting in agents. RECALL zhang2023replayenhanced focuses on improving plasticity and stability in continuous reinforcement learning through multi-head neural network training, coupled with adaptive normalization and policy distillation techniques. IQ zhang2023catastrophic employs a context partitioning strategy based on online clustering, combined with multi-head networks and knowledge distillation technology, to reduce interference between different state distributions. DRAGO fuknowledge leverages synthetic experience replay and exploration-based memory recovery to retain knowledge across tasks, mitigating catastrophic forgetting without requiring the storage of past task data.
@@
Architecture-based Methods. These methods ullah2021machine ; malagonself ; ben2022lifelong ; rusu2016progressive ; mendez2022modular concentrate on learning a policy with a set of shared parameters to handle all incremental tasks, including parameter allocation, model reorganization, and modular networks. For example, MaskNet ben2022lifelong employs a learnable modulation mask to isolate the parameters of different tasks on a fixed neural network and accelerates learning new tasks by linearly combining masks from previous tasks. CompoNet malagonself uses a modular architecture with self-composing policies to enable efficient knowledge transfer and scalable learning in continual reinforcement learning. COMP mendez2022modular decomposes complex tasks into multiple subtasks corresponding to different neural modules. These modules can be combined to form a complete policy. REWIRE NEURIPS2023_599221d7 redefines the connection methods of the neural network and reorders the neurons in each layer to achieve additional plasticity in unstable environments.
@@
Meta-Learning based Methods. These methods schmidhuber1998reinforcement ; finn2017model ; luo2022adapt ; clavera2018model assist agents in rapidly adapting to new tasks by simulating the reasoning phase during training. MB-MPO clavera2018model is a model-based method that meta-learns policies robust to ensemble dynamics, reducing reliance on accurate models. MAML finn2017model is a model-agnostic meta-learning algorithm that explicitly trains models to achieve strong generalization from a small number of gradient updates. ESCP luo2022adapt enhances the robustness and responsiveness of context encoding, significantly accelerating adaptation in reinforcement learning involving sudden environmental changes.
@@ 3 Problem Formulation
The standard Reinforcement Learning (RL) problem sutton1998reinforcement is modeled as a Markov Decision Process (MDP), with tuple $M=\langle\mathcal{S},\mathcal{A},p,r,\gamma\rangle$, where $\mathcal{S}$ is state space, $\mathcal{A}$ is action space, $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the state transition probability function, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. At each time step $t\in\mathbb{N}$, the agent observes the current state $s_{t}\in\mathcal{S}$ and takes an action $a_{t}\sim\pi(\cdot|s_{t})$, where $\pi:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the policy. The environment transitions to a new state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$, and the agent receives a reward $r(s_{t},a_{t})$. The goal of RL is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return as follows:
@@
Given $N$ tasks, the agent is expected to maximize the RL objective for each task in $\mathcal{T}=\{\tau_{1},\tau_{2},\ldots,\tau_{N}\}$. Each task $\tau_{i}$ is associated with a distinct MDP $M_{i}=\langle\mathcal{S}_{i},\mathcal{A}_{i},p_{i},r_{i},\gamma_{i}\rangle$. An RL problem is considered an instance of Continual Reinforcement Learning (CRL) if agents are required to continuously learn. Following the definition presented by Abel et al. abel2023definition , we have the following concept: an RL problem defined by a tuple $(e,\nu,\Lambda)$, where $e$ is the environment, $\nu$ is the performance function, and $\Lambda$ is the set of all agents, the problem is a CRL problem if $\forall_{\lambda^{*}\in\Lambda^{*}}\lambda^{*}\not\overset{e}{\rightsquigarrow}$ (never reaches) $\Lambda_{B}$, where $\Lambda_{B}\subset\Lambda$ is a basis such that $\Lambda_{B}\vdash_{e}$ (generates) $\Lambda$ (in $e$) and $\Lambda^{*}=\arg\max_{\lambda\in\Lambda}\nu(\lambda,e)$. In other words, the best agents continue to search indefinitely over the basis $\Lambda_{B}$ and do not converge to a fixed policy. According to our CKA-RL, whenever the number of tasks $|\mathcal{T}|$ increases, particularly when $\tau_{N+1}$ is added, a new knowledge vector $\bm{v}_{N+1}$ will be introduced. Therefore, the new task will ensure that $\lambda\not\overset{e}{\rightsquigarrow}\Lambda_{B}$, and thus continues to learn.
@@ 4.1 Knowledge Vectors
In continual reinforcement learning, leveraging knowledge from previous tasks in dynamic environments is crucial to enhance agent performance in subsequent tasks and mitigate catastrophic forgetting. Therefore, we introduce the concept of knowledge vectors to facilitate continual learning, inspired by model editing techniques ilharcoediting . Specifically, the knowledge vector, denoted as $\bm{v}_{k}\in\mathbb{R}^{d}$, represents the learned parameters that adapt the pre-trained model to a specific task $\tau_{k}$, where $d$ denotes the dimensionality of the model parameter. These vectors can be combined linearly with historical knowledge vectors to enable cross-task knowledge transfer.
@@
A linear combination of multiple knowledge vectors enables the model to reuse knowledge from previous tasks kimtest ; chen2024cross . This highlights the potential of knowledge vectors to facilitate continual knowledge adaptation, effectively tackling key challenges in continual learning.
@@ 5.1 Experimental Settings
Benchmarks. We follow the experimental settings established in prior work malagonself and compare CKA-RL with SOTA CRL methods across three distinct dynamic task sequences, including 1) Meta-World yu2020meta , 2) Freeway machado2018revisiting , and 3) SpaceInvaders machado2018revisiting . These sequences are designed to evaluate the robustness and generalization capabilities of different methods under varying levels of task complexity and action space characteristics. More details can be seen in Appendix C.
Metrics. Following standard evaluation metrics in CRL yang2023continual , we report key metrics, including average performance and forward transfer. The average performance at step $t$, denoted as $P(t)$:
@@
where $p_{i}(t)\in[0,1]$ is the success rate on task $i$ at step $t$, and each of the $N$ tasks is trained for $\Delta$ steps, where $N$ is the number of tasks, so the total number of steps is $T=N\cdot\Delta$. Its final value $P(T)$, serving as a conventional evaluation metric in CRL malagonself ; wolczyk2021continual , effectively captures the model’s stable performance across dynamic environments. The forward transfer measures the extent to which a CRL method is able to transfer knowledge across tasks. It is computed as the normalized area between the training curve of the measured run and the training curve of a reference model trained from scratch. Let $p_{i}^{b}\in[0,1]$ denote the reference performance. The forward transfer on task $i$, denoted as $FT_{i}$, is defined as follows:
@@
Methods. We compare the CKA-RL with nine SOTA methods. 1) Baseline. 2) FT-1 (Fine-Tuning Single Model) wolczyk2021continual . 3) FT-N (Fine-Tuning with Model Preservation) wolczyk2021continual . 4) ProgNet rusu2016progressive . 5) PackNet mallya2018packnet . 6) MaskNet ben2022lifelong . 7) CReLUs abbas2023loss . 8) CompoNet malagonself . 9) CbpNet dohare2024loss .
Implementation Details. We follow the prior work malagonself , employing SAC haarnoja2018soft for Meta-World and PPO schulman2017proximal for Freeway and SpaceInvaders. For high-dimensional Atari inputs (210×160 RGB), a CNN encoder maps images to compact latent features. All tasks are trained for $\Delta=1M$ steps. We use Adam (momentum $0.9$, second moment $0.999$), with batch sizes $1024/128$ and learning rates $2.5\times 10^{-4}/1\times 10^{-3}$ for PPO/SAC. The discount factor is $\gamma=0.99$. For SAC, the action standard deviation is constrained to $e^{-20},e^{2}]$, with target smoothing coefficient $5\times 10^{-3}$, auto-tuned entropy coefficient $0.2$, and action noise clipped to $0.5$. Learning starts after $5\times 10^{3}$ steps using $10^{4}$ random actions for exploration. Policy and target networks are updated every $2$ and $1$ steps, respectively, using 3-layer MLPs with 256 hidden units. For PPO, we apply GAE with $\lambda=0.95$ across 8 parallel environments, gradient clipping at $0.5$, PPO clip of $0.2$, entropy coefficient $0.01$, and 128 rollout steps. The agent uses a 2-layer MLP with 512 units, and advantage normalization is employed. Following wolczyk2022disentangling ; malagonself ; yang2023continual , CRL is applied only to the actor, while the critic is reinitialized at each task. We conduct experiments using 10 different random seeds to ensure the robustness and reliability of the results.

### v2

@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful paradigm in machine learning, enabling agents to learn optimal behaviors through interactions with dynamic environments Kaelbling et al. (1996); Dohare et al. (2024); Sun et al. (2025). RL has achieved significant success in fields such as robotic control Johannink et al. (2019); Meng et al. (2025), embodied intelligence Chen et al. (2022); Gupta et al. (2021), and natural language processing Hu et al. (2025a); Guo et al. (2025); Hu et al. (2025b). However, traditional RL usually assumes a static environment where tasks and data distributions remain fixed, intending to solve a single, well-defined problem Malagon et al. (2024). In contrast, real-world environments are typically non-stationary, requiring agents to continuously adapt to new tasks and evolving conditions. In light of this, Continual Reinforcement Learning (CRL) Khetarpal et al. (2022) has been introduced to enable agents to adapt and maintain performance across multiple tasks, facilitating more robust decision-making in dynamic environments Gaya et al. (2023); Abel et al. (2023); Wang et al. (2024).
@@
Unfortunately, CRL still faces several challenges, which are as follows. 1) Cross-task conflict: In continual learning, tasks may share certain structures or knowledge while also having incompatible goals or constraints. This interplay complicates the direct reuse of previously learned information Kumaran et al. (2016). 2) Catastrophic forgetting: When learning new tasks, the agent may overwrite or distort previously acquired knowledge, leading to a loss of performance on earlier tasks Malagon et al. (2024); Wang et al. (2024).
@@
Recently, several attempts have been proposed to enable agents to continuously learn across multiple tasks while maintaining or enhancing their performance on previously acquired tasks Zheng et al. (2025); Hu et al. (2024). The existing methods can be broadly categorized into four main groups Wang et al. (2024). The regularization-based methods Igl et al. (2021); Muppidi et al. (2024) introduce regularization terms into the learning objective that penalize large updates to important parameters for previously learned tasks. The rehearsal-based methods Daniels et al. (2022); Zhang et al. (2023a, b) mitigate forgetting by storing previous experiences in memory and periodically replaying them during the learning of new tasks. By leveraging past experiences, these methods help reduce short-term biases and improve task performance across the sequence. The architecture-based methods Ullah et al. (2021); Malagon et al. (2024); Ben-Iwhiwhu et al. (2023); Rusu et al. (2016) focus on learning a shared structure, such as modularity or composition, to facilitate continual learning. They reuse parts of previous solutions by forming abstract concepts or skills. The meta-learning based methods Schmidhuber et al. (1998); Finn et al. (2017); Luo et al. (2022); Clavera et al. (2018) improve the learning efficiency of agents by utilizing past successes and failures to refine their optimization processes. This creates an inductive bias that enhances sample efficiency and adaptability in acquiring new behaviors.
@@
Although there has been significant progress in existing methods, they still face several limitations as follows. Firstly, existing methods, such as PackNet Mallya and Lazebnik (2018), often fail to effectively address task dependencies and conflicts, which makes it challenging to transfer knowledge across tasks without interference or degradation in performance. Secondly, most methods still exhibit substantial performance degradation on previously learned tasks upon acquisition of new ones. This issue is particularly pronounced in methods that rely on regularization, as they may struggle to efficiently retain knowledge from earlier tasks as the task sequence lengthens. Lastly, many methods face scalability issues, especially as the number of tasks grows significantly, leading to increased memory and computational costs, such as CompoNet Malagon et al. (2024).
@@ 2 Related Work
Continual Reinforcement Learning (CRL) aims to enable agents to continuously learn and optimize strategies in dynamic environments, improving their ability to adapt to environmental changes and achieve goals more efficiently. The existing methods can be broadly categorized into four main types Wang et al. (2024): regularization-based methods, rehearsal-based methods, architecture-based methods, and meta-learning-based methods, each of which is detailed as:
@@
Regularization-based Methods. These methods Igl et al. (2021); Muppidi et al. (2024) add regularization terms to the training process to balance learning new and old tasks without storing models of old tasks. ITER Igl et al. (2021) periodically distills the current policy and value function into a newly initialized network during training and imitates the teacher network using a linear combination of loss terms to enhance model generalization. TRAC Muppidi et al. (2024) adaptively adjusts the regularization strength based on online convex optimization theory to prevent excessive weight drift, thereby reducing plasticity loss and enabling rapid adaptation to new distribution changes. Anand et al. Anand and Precup (2023) propose a method that decomposes the value function into permanent and transient components to address the stability-plasticity dilemma in continual reinforcement learning, enabling efficient adaptation and control in dynamic environments.
@@
Rehearsal-based Methods. These methods Daniels et al. (2022); Zhang et al. (2023a, b); Fu et al. (2025) utilize experience replay, generation replay, parameter replay, etc., to store past experiences and prevent catastrophic forgetting in agents. RECALL Zhang et al. (2023b) focuses on improving plasticity and stability in continuous reinforcement learning through multi-head neural network training, coupled with adaptive normalization and policy distillation techniques. IQ Zhang et al. (2023a) employs a context partitioning strategy based on online clustering, combined with multi-head networks and knowledge distillation technology, to reduce interference between different state distributions. DRAGO Fu et al. (2025) leverages synthetic experience replay and exploration-based memory recovery to retain knowledge across tasks, mitigating catastrophic forgetting without requiring the storage of past task data.
@@
Architecture-based Methods. These methods Ullah et al. (2021); Malagon et al. (2024); Ben-Iwhiwhu et al. (2023); Rusu et al. (2016); Mendez et al. (2022) concentrate on learning a policy with a set of shared parameters to handle all incremental tasks, including parameter allocation, model reorganization, and modular networks. For example, MaskNet Ben-Iwhiwhu et al. (2023) employs a learnable modulation mask to isolate the parameters of different tasks on a fixed neural network and accelerates learning new tasks by linearly combining masks from previous tasks. CompoNet Malagon et al. (2024) uses a modular architecture with self-composing policies to enable efficient knowledge transfer and scalable learning in continual reinforcement learning. COMP Mendez et al. (2022) decomposes complex tasks into multiple subtasks corresponding to different neural modules. These modules can be combined to form a complete policy. REWIRE Sun and Mu (2023) redefines the connection methods of the neural network and reorders the neurons in each layer to achieve additional plasticity in unstable environments.
@@
Meta-Learning based Methods. These methods Schmidhuber et al. (1998); Finn et al. (2017); Luo et al. (2022); Clavera et al. (2018) assist agents in rapidly adapting to new tasks by simulating the reasoning phase during training. MB-MPO Clavera et al. (2018) is a model-based method that meta-learns policies robust to ensemble dynamics, reducing reliance on accurate models. MAML Finn et al. (2017) is a model-agnostic meta-learning algorithm that explicitly trains models to achieve strong generalization from a small number of gradient updates. ESCP Luo et al. (2022) enhances the robustness and responsiveness of context encoding, significantly accelerating adaptation in reinforcement learning involving sudden environmental changes.
@@ 3 Problem Formulation
The standard Reinforcement Learning (RL) problem Sutton et al. (1998) is modeled as a Markov Decision Process (MDP), with tuple $M=\langle\mathcal{S},\mathcal{A},p,r,\gamma\rangle$, where $\mathcal{S}$ is state space, $\mathcal{A}$ is action space, $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the state transition probability function, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. At each time step $t\in\mathbb{N}$, the agent observes the current state $s_{t}\in\mathcal{S}$ and takes an action $a_{t}\sim\pi(\cdot|s_{t})$, where $\pi:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the policy. The environment transitions to a new state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$, and the agent receives a reward $r(s_{t},a_{t})$. The goal of RL is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return as follows:
@@
Given $N$ tasks, the agent is expected to maximize the RL objective for each task in $\mathcal{T}=\{\tau_{1},\tau_{2},\ldots,\tau_{N}\}$. Each task $\tau_{i}$ is associated with a distinct MDP $M_{i}=\langle\mathcal{S}_{i},\mathcal{A}_{i},p_{i},r_{i},\gamma_{i}\rangle$. An RL problem is considered an instance of Continual Reinforcement Learning (CRL) if agents are required to continuously learn. Following the definition presented by Abel et al. Abel et al. (2023), we have the following concept: an RL problem defined by a tuple $(e,\nu,\Lambda)$, where $e$ is the environment, $\nu$ is the performance function, and $\Lambda$ is the set of all agents, the problem is a CRL problem if $\forall_{\lambda^{*}\in\Lambda^{*}}\lambda^{*}\not\overset{e}{\rightsquigarrow}$ (never reaches) $\Lambda_{B}$, where $\Lambda_{B}\subset\Lambda$ is a basis such that $\Lambda_{B}\vdash_{e}$ (generates) $\Lambda$ (in $e$) and $\Lambda^{*}=\arg\max_{\lambda\in\Lambda}\nu(\lambda,e)$. In other words, the best agents continue to search indefinitely over the basis $\Lambda_{B}$ and do not converge to a fixed policy. According to our CKA-RL, whenever the number of tasks $|\mathcal{T}|$ increases, particularly when $\tau_{N+1}$ is added, a new knowledge vector $\bm{v}_{N+1}$ will be introduced. Therefore, the new task will ensure that $\lambda\not\overset{e}{\rightsquigarrow}\Lambda_{B}$, and thus continues to learn.
@@ 4.1 Knowledge Vectors
In continual reinforcement learning, leveraging knowledge from previous tasks in dynamic environments is crucial to enhance agent performance in subsequent tasks and mitigate catastrophic forgetting. Therefore, we introduce the concept of knowledge vectors to facilitate continual learning, inspired by model editing techniques Ilharco et al. (2023). Specifically, the knowledge vector, denoted as $\bm{v}_{k}\in\mathbb{R}^{d}$, represents the learned parameters that adapt the pre-trained model to a specific task $\tau_{k}$, where $d$ denotes the dimensionality of the model parameter. These vectors can be combined linearly with historical knowledge vectors to enable cross-task knowledge transfer.
@@
A linear combination of multiple knowledge vectors enables the model to reuse knowledge from previous tasks Kim et al. (2025); Chen et al. (2024). This highlights the potential of knowledge vectors to facilitate continual knowledge adaptation, effectively tackling key challenges in continual learning.
@@ 5.1 Experimental Settings
Benchmarks. We follow the experimental settings established in prior work Malagon et al. (2024) and compare CKA-RL with SOTA CRL methods across three distinct dynamic task sequences, including 1) Meta-World Yu et al. (2020), 2) Freeway Machado et al. (2018), and 3) SpaceInvaders Machado et al. (2018). These sequences are designed to evaluate the robustness and generalization capabilities of different methods under varying levels of task complexity and action space characteristics. More details can be seen in Appendix C.
Metrics. Following standard evaluation metrics in CRL Yang et al. (2023), we report key metrics, including average performance and forward transfer. The average performance at step $t$, denoted as $P(t)$:
@@
where $p_{i}(t)\in[0,1]$ is the success rate on task $i$ at step $t$, and each of the $N$ tasks is trained for $\Delta$ steps, where $N$ is the number of tasks, so the total number of steps is $T=N\cdot\Delta$. Its final value $P(T)$, serving as a conventional evaluation metric in CRL Malagon et al. (2024); Wołczyk et al. (2021), effectively captures the model’s stable performance across dynamic environments. The forward transfer measures the extent to which a CRL method is able to transfer knowledge across tasks. It is computed as the normalized area between the training curve of the measured run and the training curve of a reference model trained from scratch. Let $p_{i}^{b}\in[0,1]$ denote the reference performance. The forward transfer on task $i$, denoted as $FT_{i}$, is defined as follows:
@@
Methods. We compare the CKA-RL with nine SOTA methods. 1) Baseline. 2) FT-1 (Fine-Tuning Single Model) Wołczyk et al. (2021). 3) FT-N (Fine-Tuning with Model Preservation) Wołczyk et al. (2021). 4) ProgNet Rusu et al. (2016). 5) PackNet Mallya and Lazebnik (2018). 6) MaskNet Ben-Iwhiwhu et al. (2023). 7) CReLUs Abbas et al. (2023). 8) CompoNet Malagon et al. (2024). 9) CbpNet Dohare et al. (2024).
Implementation Details. We follow the prior work Malagon et al. (2024), employing SAC Haarnoja et al. (2018) for Meta-World and PPO Schulman et al. (2017) for Freeway and SpaceInvaders. For high-dimensional Atari inputs (210×160 RGB), a CNN encoder maps images to compact latent features. All tasks are trained for $\Delta=1M$ steps. We use Adam (momentum $0.9$, second moment $0.999$), with batch sizes $1024/128$ and learning rates $2.5\times 10^{-4}/1\times 10^{-3}$ for PPO/SAC. The discount factor is $\gamma=0.99$. For SAC, the action standard deviation is constrained to $e^{-20},e^{2}]$, with target smoothing coefficient $5\times 10^{-3}$, auto-tuned entropy coefficient $0.2$, and action noise clipped to $0.5$. Learning starts after $5\times 10^{3}$ steps using $10^{4}$ random actions for exploration. Policy and target networks are updated every $2$ and $1$ steps, respectively, using 3-layer MLPs with 256 hidden units. For PPO, we apply GAE with $\lambda=0.95$ across 8 parallel environments, gradient clipping at $0.5$, PPO clip of $0.2$, entropy coefficient $0.01$, and 128 rollout steps. The agent uses a 2-layer MLP with 512 units, and advantage normalization is employed. Following Wolczyk et al. (2022); Malagon et al. (2024); Yang et al. (2023), CRL is applied only to the actor, while the critic is reinitialized at each task. We conduct experiments using 10 different random seeds to ensure the robustness and reliability of the results.